OpenStudy (anonymous):

hmmm .. this is the problem .. claire spent 3/8 of her money and found she had P24.00 left. how much money did she have at first?

OpenStudy (anonymous):

let the earlier money be x

OpenStudy (anonymous):

so now given that chaire spent 3/8 of her money so she spent 3/8x money

OpenStudy (anonymous):

now given that the left money is P24.00

OpenStudy (anonymous):

Alright, since we don't know the original price we'll let it be "x". Now, let's translate the following statement. Claire spent 3/8 of her money (x) and found she had left = 24.00 now let's re-write it and you'll get: 3/8(x) = 24 solve for x and you'll find how much she had ^_^

OpenStudy (anonymous):

so x-(3/8)*x = 5x/8

OpenStudy (anonymous):

sorry 3x/8 = P24.00

OpenStudy (anonymous):

you're supposed to get x = 64 ^_^ give it a try

OpenStudy (anonymous):

SO 3x = P24.00 * 8

OpenStudy (anonymous):

x = (P24.00*8)/3 = 192.00/3

OpenStudy (anonymous):

so x = P64.00

OpenStudy (anonymous):

did you get it gel? :)